# Sublinear bilipschitz equivalence and the quasiisometric classification of solvable Lie groups

### v2

@@ Abstract.
We prove a product theorem for sublinear bilipschitz equivalences which generalizes the classical work of Kapovich, Kleiner and Leeb on quasiisometries between product spaces. We employ our product theorem to distinguish up to quasiisometry certain families among these groups which share the same dimension, cone-dimension and Dehn function; actually we do this by distinguishing them up to sublinear bilipschitz equivalence, which is slightly stronger. Especially, we recover the fact, recently obtained by Bourdon and Rémy with different groups, that there exists uncountably many quasiisometry classes of indecomposable, non-unimodular, higher rank solvable Lie groups.
@@ 1. Introduction
Our purpose here is to demonstrate that the reduction procedure, in the present case at the level of $\rho_{1}$, has applications to the quasiisometric classification in the class of Lie groups of exponential growth, which is disjoint from that of nilpotent groups. This stems from the following fact, relying on [Cor11]. Since the group $\rho_{1}(G)$ in the class $(\mathcal{C}_{1})$ have less complicated structure than $G$, it can happen that it splits in a direct product while $G$ did not. Under certain assumptions on the factors, we can then apply a suitable generalization of the Kapovich-Kleiner-Leeb theorems on quasiisometries between products [KKL98], in order to rule out the existence of an $O(u)$-equivalence (hence of a quasiisometry) between $G$ and $G^{\prime}$ when $\rho_{1}(G)$ and $\rho_{1}(G^{\prime})$ split as non-isomorphic direct products. This is the strategy of Theorem A and Theorem C below.
1.A. Sublinear bilipschitz equivalence and products
For a pointed metric space $(X,x_{0})$ we denote by $|x|=d(x,x_{0})$ the size of $x\in X$. A sublinear function is any real function $u\colon\mathbf{R}_{\geqslant 0}\to\mathbf{R}_{\geqslant 0}$ such that $\lim_{r\to\infty}\frac{u(r)}{r}=0$. We often write $u(x)$ instead of $u(|x|)$; this should not cause any confusion. Given two maps $f$ and $g$ between pointed metric spaces and a sublinear function $u$, we say that $f$ and $g$ are $O(u)$-close if $d\big{(}f(x),g(x)\big{)}=O\big{(}u(x)\big{)}$.
@@ 1.A. Sublinear bilipschitz equivalence and products
Let $n$ be a positive integer. Let $(X_{i})$ and $(Y_{i})$ be families of pointed metric spaces, $i=1,\ldots,n$. We say that the diagram
$\prod_{i=1}^{n}X_{i}$$\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{i}$$\scriptstyle{\phi}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\phi_{i}}$
commutes up to sublinear error if there exists a sublinear function $u$ such that $\phi_{i}\circ\pi_{i}$ and $\pi_{i}\circ\phi$ are $O(u)$-close for all $i$.
The precise definitions for the following theorem will be given in Section 2. Key examples of spaces of coarse type I are simply connected Riemannian manifolds with sectional curvature bounded above by a negative constant; examples of spaces of coarse type II are irreducible Riemannian symmetric spaces of noncompact type and of higher rank.
@@
Let $X=M\times\prod_{i=1}^{n}X_{i}$ and $Y=N\times\prod_{i=1}^{m}Y_{j}$ be product metric spaces. Assume the metric spaces $X_{i}$ and $Y_{j}$ are of coarse type I or II in the sense of Kapovich, Kleiner and Leeb [KKL98], and $M$ and $N$ are geodesic metric spaces with asymptotic cones homeomorphic to $\mathbf{R}^{p}$ and $\mathbf{R}^{q}$ respectively. Let $u$ be a subadditive sublinear function, and let $\phi\colon X\to Y$ be an $O(u)$-bilipschitz equivalence. Then $p=q$, $n=m$, and there exists a bijection $\sigma:\{1,\ldots,n\}\to\{1,\ldots,n\}$ and, for every $i\in\{1,\ldots,n\}$, an $O(u)$-bilipschitz equivalence $\phi_{i}:X_{i}\to Y_{\sigma(i)}$ such that the diagram
$M\times\prod_{i=1}^{n}X_{i}$$N\times\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{\sigma(i)}$$\scriptstyle{\phi}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{\sigma(i)}}$$\scriptstyle{\phi_{i}}$
commutes up to sublinear error. Moreover, this error is in the class $O(u)$.
@@
In the statement above, we do not explicitly specify how the distance on $X$ and $Y$ is built from that of the factors. In the proof we work with the $\ell^{2}$ (or Pythagorean product) metric, which is more natural when the factors are Riemannian. However, the theorem applies if one considers any distances on $X$ and $Y$ quasiisometric to the $\ell^{2}$ distances, e.g. the $\ell^{1}$ distance.
@@
When $p$ or $q$ vanish, the conclusion of Theorem A can be stated in the following way: $\phi$ is $O(u)$-close to the composition of a map of the form
$(x_{1},\ldots,x_{n})\mapsto(\phi_{1}(x_{1}),\ldots,\phi_{n}(x_{n})),$
where the $\phi_{i}$ are $O(u)$-bilipschitz equivalences, and a map permuting the factors.
Theorem A generalizes [KKL98, Theorem B], which is the case $u=1$. We state a first application below, which uses some of the results of [Pal20a, Gra23]. By pluriisometry between symmetric spaces we mean that we allow a rescaling in each factor of the de Rham decomposition.
Corollary B.
Let $X$ and $Y$ be two Riemannian globally symmetric spaces with no compact factors. Let $u$ be a subadditive sublinear function. If $X$ and $Y$ are $O(u)$-bilipschitz equivalent, then $X$ and $Y$ are pluriisometric.
Before stating the second application below, we need to recall Cornulier’s $\rho_{1}$ reduction (see §3.A for a more comprehensive account). Given a completely solvable group $S$ we denote by $\operatorname{R}_{\exp}S$ the smallest normal subgroup such that $S/\operatorname{R_{\exp}}S$ is nilpotent; this is also the intersection of the descending central series of $S$. The homomorphism $\alpha\colon S\to\operatorname{Aut}(\operatorname{R}_{\exp}S)=\operatorname{Aut}(\operatorname{Lie}(\operatorname{R}_{\exp}S))$ determined by the Adjoint action of $S$ is algebraic over $\mathbf{R}$, and can be decomposed into $\alpha=\alpha_{\sigma}\alpha_{\nu}$, where $\alpha_{\sigma}$ is valued in a diagonal $\mathbf{R}$-torus of $\operatorname{Aut}(\operatorname{Lie}(\operatorname{R}_{\exp}S))$, and $\alpha_{\nu}$ is valued in the unipotent radical of $\operatorname{Aut}(\operatorname{Lie}(\operatorname{R}_{\exp}S))$. Since $\operatorname{R}_{\exp}S$ is nilpotent (it is contained in $[S,S]$), it sits in $\ker\alpha_{\sigma}$, so that $\alpha_{\sigma}$ defines a homomorphism $S/\operatorname{R}_{\exp}S\to\operatorname{Aut}(\operatorname{R}_{\exp}S)$, that we still denote $\alpha_{\sigma}$.
@@
Theorem C.
Let $S$ and $S^{\prime}$ be two quasiisometric completely solvable groups. Assume that
@@
A more sophisticated (and slightly more general) version of Theorem C will be given in Theorem 3.16; in particular, the geometric assumption on curvature in (2) can be reformulated in terms of the structure of $H$ with the help of Heintze’s theorem [Hei74].
Theorems A and C allow us to distinguish between several families of completely solvable groups up to quasiisometry. The following is an example of this strategy in dimension $4$.
@@
Let $\alpha\in(0,1)$. The groups $S=G^{0}_{4,9}$ and $S^{\prime}=\mathbf{R}\times G_{3,5}^{\alpha}$ (the names are from the classification in [Mub63]) are the four dimensional, completely solvable Lie groups whose respective Lie algebras $\mathfrak{g}_{4,9}^{0}$ and $\mathbf{R}\times\mathfrak{g}_{3,5}^{\alpha}$ are spanned by $e_{1},\ldots,e_{4}$, subject to the following nonzero brackets:
@@
It follows from Theorem C that $S$ and $S^{\prime}$ are not quasiisometric. Here, $\rho_{1}(S)$ splits as a direct product, that is $\mathbf{R}\times P$ where $P$ is the maximal completely solvable subgroup of $\mathrm{SO}(3,1)$. See section 3.B for a detailed proof of this example.
@@
For completely solvable groups, primary quasiisometry invariants are the cone dimension (that is, the covering dimension of the asymptotic cone) and the Dehn function. In a subsequent paper [GP], we use the wide range of tools developed in [CT17] by Cornulier and Tessera in order to determine the Dehn function of all completely solvable groups of exponential growth and dimension up to 5. In Corollary 4.2 below we will summarize the contribution of our strategy to the quasiisometric classification of these groups. More precisely, Corollary 4.2 lists the $5$-dimensional groups which share the same cone dimension and Dehn function, but for which Theorem C implies that they are nonetheless quasiisometrically distinct. We recover a fact already obtained recently (in a more specific form that we explain below) by Bourdon and Rémy [BR23].
Corollary D.
@@
In order to formulate our final question, let us call a group of type NPC a completely solvable group with a CAT(0) left-invariant proper Riemannian metric, as characterized in [AW76]. Examples of groups of type NPC are provided by the $AN$ subgroups of (higher rank) semisimple Lie groups $G=KAN$, but this family is considerably more vast; in dimension $5$ already, it includes the uncountable family studied by Bourdon and Rémy and mentionned above.
Question 1.9.
In Theorem A, can one allow the factors $Y_{j}$ to be taken among non-abelian, indecomposable groups of type NPC?
The question can equally be raised in the weaker form, for the original version of the Kapovich-Kleiner-Leeb product theorem instead of Theorem A. Among all the examples in dimension $5$ that we reviewed in Section 4, none of Cornulier’s $\rho_{1}$-reductions (which are $O(\log)$-bilipschitz equivalences) could rule out a positive answer to Question 1.9.

### v3

@@ Abstract.
We prove a product theorem for sublinear bilipschitz equivalences which generalizes the classical work of Kapovich, Kleiner and Leeb on quasiisometries between product spaces. We employ our product theorem to distinguish up to quasiisometry certain families of solvable groups which share the same dimension, cone-dimension and Dehn function; actually we do this by distinguishing them up to sublinear bilipschitz equivalence, which is slightly stronger. As an application, we recover the fact, recently obtained by Bourdon and Rémy with different groups, that there exists uncountably many quasiisometry classes of indecomposable, non-unimodular, high rank solvable Lie groups.
@@ 1. Introduction
Our purpose here is to demonstrate that the reduction procedure, in the present case at the level of $\rho_{1}$, has applications to the quasiisometric classification in the class of Lie groups of exponential growth, which is disjoint from that of nilpotent groups. This stems from the following fact, relying on [COR11]: since the group $\rho_{1}(G)$ in the class $(\mathcal{C}_{1})$ have less complicated structure than $G$, it can happen that it splits in a direct product while $G$ did not.
1.A. Main results
Given a map between product spaces $f:X_{1}\times X_{2}\rightarrow Y_{1}\times Y_{2}$, when does it descend to a map between the factors? In [KKL98], Kapovich, Kleiner and Leeb provide topological conditions on the factors to ensure that this happens for $f$ a homeomorphism. They then deduce similar statements for quasi-isometries, i.e. $O(1)$-bilipschitz equivalences, assuring that quasiisometries preserve the product structure up to a bounded error. Our main contribution is to show that, in a similar setting, $O(u)$-bilipschitz equivalence also descend to the factors and that the product structure is preserved up to a sublinear error term.
For a pointed metric space $(X,x_{0})$ we denote by $|x|=d(x,x_{0})$ the size of $x\in X$. A sublinear function is any real function $u\colon\mathbf{R}_{\geqslant 0}\to\mathbf{R}_{\geqslant 0}$ such that $\lim_{r\to\infty}\frac{u(r)}{r}=0$. Two maps $f$ and $g$ between pointed metric spaces are $O(u)$-close if there is $u_{0}\in O(u)$ such that $d\big(f(x),g(x)\big)\leq u_{0}(|x|)$, and sublinearly close if there exists such a $u$ that is sublinear.
@@ 1.A. Main results
Let $n$ be a positive integer. Let $(X_{i})$ and $(Y_{i})$ be families of pointed metric spaces, $i=1,\ldots,n$. Given $i\in\{1,\ldots,n\}$ we say that the diagram
(1.1) $\prod_{i=1}^{n}X_{i}$$\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{i}$$\scriptstyle{\phi}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\phi_{i}}$
commutes up to sublinear error if there exists a sublinear function $v$ such that $\phi_{i}\circ\pi_{i}$ and $\pi_{i}\circ\phi$ are $O(v)$-close. For $x=(x_{1},\dots,x_{n})$, we denote the product map
$\Phi(x)=\left(\prod_{i=1}^{n}\phi_{i}\right)(x):=\big(\phi_{1}(x_{1}),\dots,\phi_{n}(x_{n})\big).$
When the diagrams (1.1) commute up to error $v$ for all $i$, $\phi$ and $\Phi$ are $O(v)$-close.
We refer to Section 2 for the definitions of certain terms in the statement below. Key examples of metric spaces of coarse type I are simply connected Riemannian manifolds with sectional curvature bounded above by a negative constant; examples of spaces of coarse type II are irreducible Riemannian symmetric spaces of noncompact type and high rank.
@@
Let $X=M\times\prod_{i=1}^{n}X_{i}$ and $Y=N\times\prod_{i=1}^{m}Y_{j}$, where $X_{i}$ and $Y_{j}$ are of coarse type I or II in the sense of Kapovich, Kleiner and Leeb [KKL98], and $M$ and $N$ are geodesic spaces with asymptotic cones homeomorphic to $\mathbf{R}^{p}$ and $\mathbf{R}^{q}$ respectively. Let $u$ be a subadditive nondecreasing sublinear function, and let $\phi\colon X\to Y$ be an $O(u)$-bilipschitz equivalence. Then $p=q$, $n=m$, and there is an $O(u)$-bilipschitz equivalence $\overline{\phi}:\overline{X}=\prod X_{i}\rightarrow\prod Y_{i}=\overline{Y}$ such that the maps $\overline{\pi}\circ\phi$ and $\overline{\phi}\circ\overline{\pi}$ are $O(v)$-close for some subadditive nondecreasing sublinear function $\overline{v}$.
Moreover, there exists a bijection $\sigma:\{1,\ldots,n\}\to\{1,\ldots,n\}$ and, for every $i\in\{1,\ldots,n\}$, an $O(u)$-bilipschitz equivalence $\phi_{i}:X_{i}\to Y_{\sigma(i)}$ such that the diagram
$M\times\prod_{i=1}^{n}X_{i}$$N\times\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{\sigma(i)}$$\scriptstyle{\overline{\phi}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{\sigma(i)}}$$\scriptstyle{\phi_{i}}$
commutes up to a sublinear error $v^{\prime}$ that is subadditive and nondecreasing.
The case $u=1$ follows from [KKL98, Theorem B]; in this case $v$ and $v^{\prime}$ can be shown to be bounded (see Remark 1.6); a more elementary proof in the case $n=1$ and $X_{1}$ and $Y_{1}$ are non-elementary regular trees was given in [SW02].
We state a first application below, which uses some of the results of [PAL20a, GRA23]. By pluriisometry between symmetric spaces we mean that we allow a rescaling in each factor of the de Rham decomposition.
Corollary B.
Let $X$ and $Y$ be two Riemannian globally symmetric spaces with no compact factors. If $X$ and $Y$ are sublinear bilipschitz equivalent, then $X$ and $Y$ are pluriisometric.
Corollary B restricted to quasiisometries instead of sublinear bilispchitz equivalences is due to Kleiner-Leeb [KL97, Corollary 1.1.4] and builds on the work of Mostow [MOS71] for the part concerning the rank one factors.
The proof of Theorem A relies on Theorem C, which is the main innovation in this paper.
Definition 1.3.
A homeomorphism $F:\prod_{i=1}^{n}\widetilde{X}_{i}\rightarrow\prod_{i=1}^{m}\widetilde{Y}_{i}$ between a product of metric spaces is said to preserve the product structure if up to reindexing the factors there are homeomorphisms $F_{i}$ such that for each $i$, the following diagram commutes:
$\prod_{i=1}^{n}\widetilde{X}_{i}$$\prod_{i=1}^{n}\widetilde{Y}_{i}$$\widetilde{X}_{i}$$\widetilde{Y}_{i}.$$\scriptstyle{F}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{i}}$$\scriptstyle{F_{i}}$
Theorem C.
Let $X_{i}$ and $Y_{j}$ be geodesic metric spaces, and $u$ be a subadditive nondecreasing sublinear function. Suppose $\phi:X=\prod_{i=1}^{n}X_{i}\rightarrow\prod_{i=1}^{n}Y_{j}=Y$ is an $O(u)$-bilipschitz equivalence such that for all $u$-admissible cones $X_{\omega}$ and $Y_{\omega}$, the cone map $\operatorname{Cone}(\phi):X_{\omega}\rightarrow Y_{\omega}$ preserves the product structure. Then there exists a bijection $\sigma:\{1,\ldots,n\}\to\{1,\ldots,n\}$ and, for every $i\in\{1,\ldots,n\}$, an $O(u)$-bilipschitz equivalence $\phi_{i}:X_{i}\to Y_{\sigma(i)}$ such that the diagram
$\prod_{i=1}^{n}X_{i}$$\prod_{i=1}^{n}Y_{i}$$X_{i}$$Y_{\sigma(i)}$$\scriptstyle{\phi}$$\scriptstyle{\pi_{i}}$$\scriptstyle{\pi_{\sigma(i)}}$$\scriptstyle{\phi_{i}}$
commutes up to sublinear error $v$, that is subadditive and nondecreasing. Especially, $\phi$ is $O(v)$-close to the product map $\Phi=\prod_{i=1}^{n}\phi_{i}$ of Definition 1.2.
Most of Section 2 is devoted to the proof of Theorem C. The other ingredient needed for Theorem A is the topological splitting result from [KKL98], stating that the spaces considered in Theorem A satisfy the hypothesis of Theorem C.
Theorem 1.4 ([KKL98], Theorem 5.1).
Suppose $X_{i}$ and $Y_{j}$ are spaces of coarse type I or II. Denote $\overline{X}:=\prod_{i=1}^{n}X_{i},\overline{Y}:=\prod_{i=1}^{n}Y_{i}$, and let $\overline{X}_{\omega},\overline{Y}_{\omega}$ be respective asymptotic cones. Suppose $F:=\mathbf{R}^{p}\times\overline{X}_{\omega}\rightarrow\mathbf{R}^{q}\times\overline{Y}_{\omega}$ is a homeomorphism. Then $p=q$, $n=m$, and the homeomorphism $\overline{F}:\overline{X}_{\omega}\rightarrow\overline{Y}_{\omega}$ given by $\pi_{\overline{Y}_{\omega}}\circ F_{\restriction\overline{X}_{\omega}}$ preserves the product structure.
Theorem A is proved in Section 2.D. In addition to Theorem C and Theorem 1.4, it uses Proposition 2.22, which allows one to deal with Euclidean factors.
@@
In the statements above, we do not explicitly specify how the distance on the product is built from that of the factors. In the proof we work with the $\ell^{2}$ (or Pythagorean product) metric, which is more natural when the factors are Riemannian. However, the theorem applies if one considers any distances on $X$ and $Y$ quasiisometric to the $\ell^{2}$ distances, e.g. the $\ell^{1}$ distance.
The following remarks concern the two differences between our product theorem and the one for quasiisometries in [KKL98].
@@
When $\phi$ is a quasiisometry, i.e. an $O(1)$-equivalence, the diagrams in the statements of Theorems A and C commute up to a bounded error $D\in O(1)$ ([KKL98], Theorem A). For $O(u)$-equivalences we can only bound this error term by an arbitrary sublinear function $v$ that may not be in $O(u)$. We use Lemmas 2.1 and 2.2 to upgrade $v$ to a nondecreasing subadditive function, which are natural assumptions in the context of sublinear bilipschitz equivalence [PAL18, GRA23]. We are not aware of any application that may be affected by the fact that $v\notin O(u)$, not even when $u=1$ and $\phi$ is a quasiisometry. We stress that nonetheless, the factor maps $\phi_{i}:X_{i}\rightarrow Y_{i}$ are $O(u)$-equivalences.
Remark 1.7.
One notable and delicate difference between our statements those of [KKL98] concern the nontranslatability condition: a pair of metric spaces $(Z,W)$ is nontranslatable if any two homeomorphisms $f,g:Z\rightarrow W$ that are at finite distance must coincide ([KKL98, Definitions 2.2, 2.3]). In [KKL98], this notion is used in two separate ways. First, it is one of the key definitions in the context of their topological splitting (Theorem 1.4 above). The second use of nontranslatability in [KKL98] is as a means to upgrade the error term in the commuting diagrams of Theorems A and C to an $O(1)$ error.
In our work we do not require this notion of nontranslatability outside of Theorem 1.4, which may be taken as a black box in this paper. While the assumptions of Theorem C must be strengthened in order to ensure that the error term is in $O(u)$ (see Example 1.8), we could not determine whether nontranslatability of the pairs of factors would be a sufficient assumption.
Example 1.8.
Let $\mathbf{R}_{1},\mathbf{R}_{2}$ be two copies of the line $\mathbf{R}$ (a non non-translatable space). Consider the map $\phi:\mathbf{R}_{1}\times\mathbf{R}_{2}\rightarrow\mathbf{R}_{1}\times\mathbf{R}_{2}$ given by $(x,y)\mapsto(x+y^{\frac{1}{3}},y+x^{\frac{1}{3}})$. One may check it is a quasiisometry, and that in any asymptotic cone, the cone map preserves the product structure. The map $\phi_{1}:\mathbf{R}_{1}\rightarrow\mathbf{R}_{1}$ defined as $\phi_{1}(x)=\pi_{1}\circ\phi(x,0)$ is a quasiisometry. However, for all $y\in\mathbf{R}$
$d(\phi_{1}\circ\pi_{1}(x,y),\pi_{1}\circ\phi(x,y))=|y|^{\frac{1}{3}}$
which is not a bounded function of $|y|$.
Before stating the second application of Theorem A, we need to recall Cornulier’s $\rho_{1}$ reduction (see §3.A for a more comprehensive account). Given a completely solvable group $S$ we denote by $\operatorname{R}_{\exp}S$ the smallest normal subgroup such that $S/\operatorname{R_{\exp}}S$ is nilpotent; this is also the intersection of the descending central series of $S$. The homomorphism $\alpha\colon S\to\operatorname{Aut}(\operatorname{R}_{\exp}S)=\operatorname{Aut}(\operatorname{Lie}(\operatorname{R}_{\exp}S))$ determined by the Adjoint action of $S$ is algebraic over $\mathbf{R}$, and can be decomposed into $\alpha=\alpha_{\sigma}\alpha_{\nu}$, where $\alpha_{\sigma}$ is valued in a diagonal $\mathbf{R}$-torus of $\operatorname{Aut}(\operatorname{Lie}(\operatorname{R}_{\exp}S))$, and $\alpha_{\nu}$ is valued in the unipotent radical of $\operatorname{Aut}(\operatorname{Lie}(\operatorname{R}_{\exp}S))$. Since $\operatorname{R}_{\exp}S$ is nilpotent (it is contained in $[S,S]$), it sits in $\ker\alpha_{\sigma}$, so that $\alpha_{\sigma}$ defines a homomorphism $S/\operatorname{R}_{\exp}S\to\operatorname{Aut}(\operatorname{R}_{\exp}S)$, that we still denote $\alpha_{\sigma}$.
@@
Theorem D.
Let $S$ and $S^{\prime}$ be two completely solvable groups. Assume that there exists a sublinear, subadditive function $u$ such that $S$ and $S^{\prime}$ are $O(u)$-sublinear bilipschitz equivalent (which holds, in particular, if $S$ and $S^{\prime}$ are quasiisometric with $u=1$ in that case). Further assume that
@@
A more sophisticated (and slightly more general) version of Theorem D will be given in Theorem 3.16; in particular, the geometric assumption on curvature in (2) can be reformulated in terms of the structure of $H$ with the help of Heintze’s theorem [HEI74].
Theorem A and Theorem D allow us to distinguish between several families of completely solvable groups up to quasiisometry. The following is an example of this strategy in dimension $4$.
@@
Let $\alpha\in(0,1)$. The groups $S=G^{0}_{4,9}$ and $S^{\prime}_{\alpha}=\mathbf{R}\times G_{3,5}^{\alpha}$ (the names are from the classification in [MUB63]) are the four dimensional, completely solvable Lie groups whose respective Lie algebras $\mathfrak{g}_{4,9}^{0}$ and $\mathbf{R}\times\mathfrak{g}_{3,5}^{\alpha}$ are spanned by $e_{1},\ldots,e_{4}$, subject to the following nonzero brackets:
@@
It follows from Theorem D (with $u\equiv 1$) that for every $\alpha\in(0,1)$, $S$ and $S^{\prime}_{\alpha}$ are not quasiisometric. Here, $\rho_{1}(S)$ splits as a direct product, that is $\mathbf{R}\times P$ where $P$ is the maximal completely solvable subgroup of $\mathrm{SO}(3,1)$. In this example the original Kapovich-Kleiner-Leeb theorem cannot be applied since $S$ is not a direct product of non-trivial Lie groups. See section 3.B for a direct proof of this example, that we use as a warm-up for the general case.
@@
For completely solvable groups, primary quasiisometry invariants are the cone dimension (that is, the covering dimension of the asymptotic cone) and the Dehn function. In a subsequent paper [GP25], we use the wide range of tools developed in [CT17] by Cornulier and Tessera in order to determine the Dehn function of all completely solvable groups of exponential growth and dimension up to 5. In Corollary 4.2 below we will summarize the contribution of our strategy to the quasiisometric classification of these groups. More precisely, Corollary 4.2 lists the $5$-dimensional groups which share the same cone dimension and Dehn function, but for which Theorem D implies that they are nonetheless quasiisometrically distinct. We recover a fact already obtained recently (in a more specific form that we explain below) by Bourdon and Rémy [BR23].
Corollary E.
@@
Bourdon and Rémy obtain their result through the determination of a critical exponent for $L^{p}$-cohomology in degree $2$; by using [PAL20b], our proof also ultimately involves (although not in the present paper) the use of a critical exponent. The latter appears at first in [PAL20b] as a variant of Pansu’s conformal dimension; however it coincides with the critical exponent for $L^{p}$-cohomology in degree one of a family Heintze groups occuring as direct factors of $\rho_{1}(G_{5,19}^{1,\beta})$ (See [BK13], [CAR17] and [PAL20b, §3.2.1]). It would thus be interesting to investigate the $L^{p}$-cohomology of $G_{5,19}^{1,\beta}$ in degree $2$.
In order to formulate our final question, let us call a group of type NPC a completely solvable group with a CAT(0) left-invariant proper Riemannian metric, as characterized in [AW76]. Examples of groups of type NPC are provided by the $AN$ subgroups of (high rank) semisimple Lie groups $G=KAN$, but this family is considerably more vast; in dimension $5$ already, it includes the uncountable family studied by Bourdon and Rémy and mentionned above.
Question 1.13.
In Theorem A, can one allow the factors to be taken among non-abelian, indecomposable groups of type NPC?
The question can equally be raised in the weaker form, for the original version of the Kapovich-Kleiner-Leeb product theorem instead of Theorem A.
Topological splitting, Theorem 1.4 [KKL98] Theorem C $\rho_{1}$– reduction, Theorem 1.10 [COR11] Theorem A Theorem 3.16 Theorem D SBE classification of symmetric spaces, Corollary B Corollary E Applications to solvable Lie groups with [PAL18] (rank one) and [GRA23] (no rank one)
Figure 1. Logical dependencies between the main statements in this paper.